3x+5/9+2x-1/9=7/9
Could you show your steps please, thanks!

Answer by rfer(16322)

5x+4/9=7/9
5x=7/9-4/9
5x=3/9
x=(3/9)/5
x=3/9*1/5
x=3/45
x=1/15